OpenStudy (anonymous):

How to determine whether the graphs of the two equations are parallel lines, perpendicular lines or neither? Example y=x+4 y=x-3

OpenStudy (anonymous):

Firstly, line equations are of the form \(y=mx+c\) Parallel lines have the same slope (i.e. \(m_1 = m_2\)) Perpendicular lines have inverse negative slopes (i.e. \(m_1 = -\frac{1}{m_2}\))

OpenStudy (anonymous):

Kindly give example on this or kindly show the solution please

OpenStudy (anonymous):

Take for example, the line \(y=3x+5\). The slope of this line is 3 (the coefficient of x). A line parallel to this will have the same slope, i.e. be in the form \(y = 3x + c\). Note that they have the same x coefficient. A line parallel to this will have the inverse negative coefficient, i.e. be in the form \(y = -\frac{1}{3}x + c\). Note that the coefficient of x is the negative inverse of 3 (i.e. \(-\frac{1}{3}\)).

OpenStudy (anonymous):

y=x+4 y=x-3 Is this parallel lines, perpendicular lines or neither? kindly show the solutions and explanation on this please

OpenStudy (anonymous):

What is the coefficient of x in both of those cases?

OpenStudy (anonymous):

no coefficient

OpenStudy (anonymous):

What do you get if you add x + x? What if you add 2x + x? 3x + x?